OpenStudy (anonymous):

find the equations of three tangent lines to the graph of y = 4x^3 -40x^2 + 125x - 126 which pass through point (1, -1) (but not on the curve)

OpenStudy (anonymous):

\[y = 4x^3 -40x^2 + 125x -126\]

OpenStudy (allank):

Cool question. I'd start by calculating the first derivative of y which will give me the slope of the lines we are looking for.

OpenStudy (allank):

Can you take it from here?

OpenStudy (anonymous):

yea

OpenStudy (allank):

Great. Good luck. :)

OpenStudy (anonymous):

thanks

OpenStudy (allank):

You're welcome :)

OpenStudy (anonymous):

i get \[y' = 12x^2 - 80x +125\] what now

OpenStudy (anonymous):

im supposed to get 3 slopes

OpenStudy (anonymous):

factor right?

OpenStudy (anonymous):

\[y = (2x-5)(6x-25)\]

OpenStudy (anonymous):

y prime i mean

OpenStudy (anonymous):

slopes so far are 5/2, 25/6, and...?

OpenStudy (allank):

Just a sec iop360...

OpenStudy (anonymous):

actually is that even right?

OpenStudy (anonymous):

ok

OpenStudy (allank):

Now, what we want to do is first get the positions where the tangents touch the curve...

OpenStudy (allank):

Think about it...how would we do that?

OpenStudy (allank):

Hint: Simultaneous equations.

OpenStudy (anonymous):

so use m = m with one m representing the first derivative i just calculated and the other m representing rise/run

OpenStudy (anonymous):

equate the two and solve for x?

OpenStudy (allank):

Sorry for taking forever to reply...

OpenStudy (anonymous):

\[\frac{ (4x^3 - 40x^2 +125x -126) - (-1) }{ x-1 }= 12x^2 -80x + 125\]

OpenStudy (anonymous):

its alright

OpenStudy (allank):

Okay, we have the gradient: m = 12x^2-80x+125 and we have a point (1,-1). I'm thinking we can use the general form of an equation of a line: y-y1 = m(x-x1) where y1 and x1 are -1 and 1 respectively, and m is the gradient. I'm not sure whether that will work. Lemme try....

OpenStudy (anonymous):

we would have to find m first

OpenStudy (allank):

True. Trying something different...

OpenStudy (anonymous):

i think i found the slopes, tell me what you have too slope = 0, 5/2, 4

OpenStudy (allank):

Well, I have used a different method, where I first determine the positions where the tangents touch the curve. Want to finish off and compare answers?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you mean now find the line equations?

OpenStudy (allank):

Yep.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

y = -1 y = (5/2)x - (7/2) y = 4x - 5

OpenStudy (anonymous):

doesnt seem right when i graph it...

OpenStudy (anonymous):

hopefully u have a better one

OpenStudy (allank):

Well, I've graphed yours and It seems okay. :) My equation is giving me imaginary solutions -_-

OpenStudy (anonymous):

u graphed mine?

OpenStudy (allank):

Yep.

OpenStudy (anonymous):

it doesnt look tangent to the curve

OpenStudy (allank):

Plotting on a calculator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

they just all intersect together and dont touch the curve

OpenStudy (allank):

Darn. Okay, maybe I'm just sleepy to notice that on a computer grapher...so wanna see my method? Though it may not lead to an answer...

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

perhaps i should bump this

OpenStudy (allank):

Okay, so we have the slope m = 12x^2-80x+125. Remember that the value of m depends on the x value on the curve.... Knowing the point (1,-1) , the equation of the tangent(s) will be y+1 = m(x-1) Plugging in m: y = m(x-1)-1 y = (12x^2-80x+125)(x-1) -1 y = 12x^3 - 80x^2 +125x - 12x^2 + 80x -125 - 1 y = 12x^3 - 92x^2 + 205x - 126 (equation 1) So, we have an equation that depends on the very same x value on the curve as the original equation: y = 4x^3 -40x^2 + 125x - 126 (equation 2) Two nice simultaneous equations that will give the x values on the curve where the tangents touch the curve: y = 12x^3 - 92x^2 + 205x - 126 (equation 1) y = 4x^3 -40x^2 + 125x - 126 (equation 2)

OpenStudy (allank):

I hope I've done the math right. Subtracting the second from the first eliminates y and the 126. y = 12x^3 - 92x^2 + 205x - 126 (equation 1) y = 4x^3 -40x^2 + 125x - 126 (equation 2) 0 = 8x^3 +52x^2 + 330x And solving for x gives x=0 and two other imaginary solutions. *sigh*

OpenStudy (anonymous):

i see

OpenStudy (allank):

Do bump @iop360. I'll sleep on it and in case you haven't gotten a valid response by the time I'm awake, I will post the solution.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

ill leave this question open

OpenStudy (anonymous):

thanks for helping

OpenStudy (anonymous):

hey, i got the question right! if you want to know how, the values for x i found (0, (5/2), and 4) were actually the x coordinates of the tangents so i plugged these values in the original equation to get their respective y values after that i used each (x,y) coordinates of the three tangents with the given point in the question (1,-1) in the rise/run formula to get the slope for each line

OpenStudy (anonymous):

plug in the found slopes in point-slope formula and it should be right

OpenStudy (allank):

Nice. Good job! :)